OpenStudy (anonymous):

The final exam is worth 20% of all yearly grade. What should student's least current grade be, so that he passes if he gets 100 on the final?

OpenStudy (anonymous):

@texaschic101 @aaronq

OpenStudy (anonymous):

passing grade is 75.

OpenStudy (mosaic):

Let the student's current grade be "x". 0.8*x + 0.2*100 = 75 Solve for x.

OpenStudy (anonymous):

\[0.8x + 0.2 \times 100 = 75\]\[0.8x + 20 = 75\]\[0.8x = 55\]\[x = 55 \div 0.8 = 68.75\]

OpenStudy (anonymous):

Right ?

OpenStudy (mosaic):

Yes!

OpenStudy (anonymous):

ty

OpenStudy (mosaic):

yw